Solve the system of equations. Enter your answer as an ordered pair.
5x - y = 27
5x + 4y = 42
Answer:

Answer by stanbon(75887)

Solve the system of equations. Enter your answer as an ordered pair.
5x - y = 27
5x + 4y = 42
-----
Subtract to get:
5y = 15
y = 3
==========
Solve for "x":
5x - y = 27
5x -3 = 27
5x = 30
x = 6
============
Cheers,
Stan H.
============